### v1

# EarthMind: Towards Multi-Granular and Multi-Sensor Earth Observation with Large Multimodal Models

Abstract
Large Multimodal Models (LMMs) have demonstrated strong performance in various vision-language tasks. However, they often struggle to comprehensively understand Earth Observation (EO) data, which is critical for monitoring the environment and the effects of human activity on it. In this work, we present EarthMind, a novel vision-language framework for multi-granular and multi-sensor EO data understanding. EarthMind features two core components: (1) Spatial Attention Prompting (SAP), which reallocates attention within the LLM to enhance pixel-level understanding; and (2) Cross-modal Fusion, which aligns heterogeneous modalities into a shared space and adaptively reweighs tokens based on their information density for effective fusion. To facilitate multi-sensor fusion evaluation, we propose EarthMind-Bench, a comprehensive benchmark with over 2,000 human-annotated multi-sensor image-question pairs, covering a wide range of perception and reasoning tasks. Extensive experiments demonstrate the effectiveness of EarthMind. It achieves state-of-the-art performance on EarthMind-Bench, surpassing GPT-4o despite being only 4B in scale. Moreover, EarthMind outperforms existing methods on multiple public EO benchmarks, showcasing its potential to handle both multi-granular and multi-sensor challenges in a unified framework.
Figure 1: The proposed EarthMind supports unified multi-granular understanding for Earth Observation (EO) imagery, including image-level, region-level, and pixel-level tasks. In addition, it enables complementary multi-sensor fusion across Optical and SAR modalities.
1 Introduction
Large multimodal models (LMMs), which integrate large language models (LLMs)  [1, 2] with visual encoders, have shown remarkable success across a variety of vision-language tasks, including image captioning [3, 4], visual question answering [5, 6, 7], and grounding [8, 9, 10]. Among various application domains, Earth Observation (EO) [11, 12, 13] is of particular importance, as it allows monitoring the Earth and the effect of human activities on it. However, LMMs trained on general-purpose images often struggle to generalize to EO data, due to a significant domain gap. Recent work has addressed this challenge by constructing large-scale instruction tuning datasets [14, 15, 16, 17, 18] specifically tailored for EO, enabling better adaptation of LMMs to this domain.
Despite recent advances, existing LMMs are limited in their understanding of EO data. Firstly, EO tasks span multiple levels of granularity, from pixel-level segmentation [19, 20], over region-level semantic understanding [21, 22], up to image-level scene classification [23, 24]. Secondly, EO data comprise multiple sensing modalities, including optical imagery (e.g., RGB and Multispectral) and Synthetic Aperture Radar (SAR) [25, 26]. These modalities are inherently complementary: optical images provide rich texture and spectral information under favorable conditions, while SAR captures structural details regardless of weather or illumination. Although various sensor types exist, effective fusion, particularly between SAR and optical data modalities, remains a key challenge for EO understanding. As summarized in Tab. 1, achieving fine-grained, multi-sensor comprehension in EO remains largely unresolved.
To tackle these challenges, we introduce EarthMind, the first LMM capable of fusing multi-sensor EO inputs and performing reasoning across multiple semantic levels, as shown in Fig. 1. It achieves this by projecting heterogeneous features from different sensors and scales into a unified semantic space, thus enabling effective interpretation by LLMs. The novelty of EarthMind lies in two key design concepts that enable spatial grounding and cross-modal understanding in EO data. First, Spatial Attention Prompting (SAP) enhances pixel-level grounding by explicitly extracting and reallocating attention to regions aligned with queried objects. This overcomes limitations of prior approaches [9, 10] that combine segmentation foundation models [27, 28] with LLMs but degrade in EO settings due to vague boundaries and scale imbalances. Second, a Cross-modal Fusion mechanism, built upon token-level contrastive learning, guides the integration of complementary modalities (e.g., RGB and SAR) into a unified semantic space. Equipped with Modality Mutual Attention, EarthMind adaptively selects the most informative features from each modality, thereby facilitating robust autoregressive learning within the LLM.
Additionally, we propose EarthMind-Bench, a new benchmark designed to evaluate LMMs in challenging EO scenarios. As shown in Tab. 1, EarthMind-Bench offers several unique features. Firstly, it encompasses multi-granular tasks, ranging from coarse-grained image understanding to fine-grained segmentation. Secondly, it introduces multi-sensor data, in particular paired RGB-SAR imagery, enabling evaluation of cross-modal fusion capabilities. Thirdly, it covers multi-level questions, spanning low-level perception as well as high-level reasoning. EarthMind-Bench consists of more than 2,000 multiple-choice and open-ended questions, providing a comprehensive benchmark to assess the ability of LMMs to interpret and reason over EO data.
We implemented EarthMind based on Qwen-2.5-3B [29], and its effectiveness is demonstrated from three key perspectives. Firstly, it achieves state-of-the-art performance on several downstream tasks, outperforming existing EO-focused LMMs of larger scale, and surpassing specialized segmentation models for the first time on pixel-level tasks. Secondly, it significantly outperforms the baselines on EarthMind-Bench across both RGB and SAR, where it improves performance through the effective fusion mechanism.
Table 1: (Left) Comparison of EarthMind with existing EO LMMs. EarthMind supports both multi-granular and multi-sensor understanding. (Right) Comparison of EarthMind-Bench with existing EO benchmarks in terms of multi-sensor support, granularity, and task level. “S” denotes single-modality input; “M” indicates multiple modalities (used independently); “F” represents paired sensor fusion. “MC” and “OE” refer to multiple-choice and open-ended formats, respectively.
Method Multi-Granular Multi-Sensor Image Region Pixel Handling Fusion RSGPT [18] GeoChat [14] EarthGPT [17] Earthmarker [30] LHRS-bot [15] SkyEyeGPT [16] Skysensegpt [16] EarthDial [31] GeoPixel [32] RSUniVLM [33] EarthMind
Benchmark Multi Sensor Multi Gran. Multi Level Task Type RSIEval [18] S MC + OE HnstD [34] S MC + OE GeoChat-Bench [14] S OE VRSBench [35] S MC + OE LHRS-Bench [15] S MC VLEO-Bench [36] S MC + OE FIT-RSRC [37] S MC UrBench [38] S MC XLRS-Bench [39] S MC + OE GEOBench-VLM [40] M MC + OE EarthMind-Bench M + F MC + OE
2 Related Work
Earth Observation LMMs. Starting from the excellent foundation of image LMMs, many works [18, 14, 16, 17, 30, 15, 37, 32, 31] try to transfer the success of general image understanding to the EO field. A key challenge is the lack of instruction-tuned EO datasets. To address this, RSGPT [18] proposed the first large-scale EO image-text paired dataset, enabling conversation tasks such as image captioning and VQA. GeoChat [14] and SkyEyeGPT [16] extend LMM capabilities to region-level visual grounding by introducing region-centric instruction data. LHRS-Bot [15] leverages large-scale available EO imagery aligned with OpenStreetMap annotations to improve VLM pretraining. To enhance complex reasoning, SkysenseGPT [37] introduces the FIT-RS dataset that focuses on understanding the relationships between spatial entities. GeoPixel [32] further pushes the boundary to pixel-level grounding by constructing an automatic pipeline to generate grounded conversations based on EO imagery. Beyond optical data, EarthDial [31] incorporates a diverse set of modalities, including Multispectral and Hyperspectral imagery and SAR, with the goal of improving generalization across heterogeneous EO data sources. Despite these advances, current EO LMMs remain limited in jointly supporting multi-granular tasks and multi-sensor fusion, which are essential for real applications.
Earth Observation Multimodal Benchmarks. The rapid progress of LMMs in the EO domain has also stimulated the development of dedicated evaluation benchmarks. RSIEval [18] provides human-annotated captions and VQA pairs to evaluate VLMs in remote sensing. LHRS-Bench [15] introduces hierarchical taxonomies to assess LMMs across multiple dimensions. VLEO-Bench [36] focuses on real-world applications, including urban monitoring, disaster relief, and land use. Beyond conversational tasks, VRSBench [35] and GeoChat-Bench [14] incorporate grounding tasks to evaluate the localization capabilities of LMMs. HnstD [34] aims to detect model hallucinations, while FIT-RSRC [37] targets reasoning about object relationships. XLRS-Bench [39] leverages ultra-high-resolution imagery for comprehensive evaluation. Concurrently, GEOBench-VLM [40] proposes a multi-task geospatial benchmark covering diverse EO tasks. Despite these advances, a notable gap remains: None of the existing benchmarks explicitly evaluates the ability of LMMs to perform multi-sensor fusion, which is a key capability in real EO scenarios.
3 EarthMind
3.1 Overview
From dense urban mapping to natural terrain analysis, EO applications require models that can unify multi-modal data to enable rich scene characterization and improved decision-making. At its core, EarthMind integrates a set of vision encoders with a language model to enable flexible and generalizable cross-modal reasoning.
As shown in Fig. 2, for spatial and semantic understanding at different levels of granularity, our model employs three specialized encoders: a visual encoder $\mathbf{E}_{v}$ for global semantic perception, a region encoder $\mathbf{E}_{r}$ for object-level understanding, and a grounding encoder $\mathbf{E}_{g}$ for fine-grained spatial segmentation. These components produce hierarchical representations of the input, which are projected into a shared language space using a vision-language projector (VLP). This yields a token sequence $X^{V}=\{x^{v}_{1},x^{v}_{2},...,x^{v}_{P}\}$ that captures both global and local visual information.
To further enable dense prediction tasks, we introduce a set of learnable segmentation tokens $X^{S}=\{x^{s}_{1},x^{s}_{2},...,x^{s}_{Q}\}$, where each token can be interpreted as a query embedding tailored to capture a specific aspect of the spatial layout. The combined sequence of visual tokens $X^{V}$, segmentation tokens $X^{S}$, and tokenized language queries is processed by an LLM, which performs joint cross-modal reasoning. The hidden states of “<SEG>” tokens and the output of $\mathbf{E}_{g}$ are passed to a lightweight mask decoder $\mathbf{D}_{m}$ to produce fine-grained segmentation maps. To further improve spatial grounding in complex EO scenes, we propose Spatial Attention Prompting (SAP), detailed in Sec. 3.2.
In parallel, our framework supports multi-sensor fusion by treating heterogeneous EO data as a form of video-like input. Inspired by recent advances in video-language models [41, 42], we adopt a unified data formatting strategy: single- and dual-channel SAR images are padded to form pseudo-RGB frames, while multispectral bands are grouped in triplets to construct temporally-ordered multi-frame sequences. These sequences mimic the temporal structure of video inputs and are processed by shared encoders, enabling our model to exploit cross-frame dependencies and spectral complementarity.
To enhance inter-modal alignment, we propose a Cross-modal Fusion module (details in Sec. 3.3) that leverages paired sensor inputs under a unified language conditioning. This module facilitates effective interaction between modalities by attending to complementary cues across spectral and spatial dimensions.
Figure 2: An overview of the EarthMind, which can handle multi-granular and multi-sensor EO data understanding. It facilitates effective interaction between modalities with Cross-modal Fusion and performs accurate pixel-level understanding with Spatial Attention Prompting.
3.2 Spatial Attention Prompting (SAP)
In the above architecture, the “<SEG>” token acts as an implicit language query that searches for potential target objects within the image. However, empirical results show a significant performance drop when applying this framework to EO scenarios. To investigate the cause of this degradation, we examine whether the “<SEG>” token effectively captures the semantics of the queried objects aligned with the query in EO imagery.
During LLM inference, tokens interact with each other through multi-head attention layers. A token that receives higher attention weights from others is considered more influential in the model’s reasoning process. Motivated by this, we measure the importance of each image token with respect to the “<SEG>” token by computing the cross-attention maps between segmentation tokens and all image tokens across all transformer layers, resulting in $A_{\text{seg}}\in\mathbb{R}^{L\times H\times Q\times P}$, where $L$ is the number of transformer layers and $H$ is the number of attention heads. We compute the attention map by applying softmax over the image token dimension, and average over the heads to obtain a Seg-to-Image attention map:
$\small\hat{A}_{\text{seg}}=\frac{1}{H}\sum_{h=1}^{H}\text{Softmax}(A_{\text{seg}}^{(l,h)}),\quad\hat{A}_{\text{seg}}\in\mathbb{R}^{L\times Q\times P}$ (1)
As shown in Fig. 2, the attention map reflects the spatial focus of the LLM’s segmentation prompts. However, EO imagery often exhibits extreme object scale variations, ambiguous textures, and weak semantic boundaries, which lead to attention drifting away from the target regions. Our attention visualizations further illustrate this phenomenon in Sec. 5.4.
To guide the model toward more semantically meaningful spatial priors, we introduce a supervision signal based on ground-truth masks. Specifically, we downsample the binary mask to match the image token resolution and treat it as a probability distribution over tokens. Then, we minimize the Kullback–Leibler (KL) divergence [43] between the normalized attention map and the target distribution:
$\small\mathcal{L}_{\text{KL}}=\sum_{l=1}^{L}\sum_{q=1}^{Q}\text{KL}\left(\hat{A}_{\text{seg}}^{(l,q)}\,\|\,G_{q}\right)$ (2)
where $G_{q}\in\mathbb{R}^{P}$ is the ground-truth token-level distribution corresponding to the $q$-th seg token. This training objective encourages the segmentation tokens to attend more accurately to the queried regions, thereby enhancing pixel-level mask prediction.
3.3 Cross-modal Fusion
Given multi-sensor input, we extract cross-modal image tokens before feeding into the LLM: $X^{V}_{o}=\{x^{o}_{1},x^{o}_{2},...,x^{o}_{P}\}$ and $X^{V}_{s}=\{x^{s}_{1},x^{s}_{2},...,x^{s}_{P}\}$, which represent features from optical (RGB) and non-optical (e.g., SAR) modalities, respectively. To fuse modalities, we propose the concepts of Modality Alignment and Modality Mutual Attention, as detailed below.
Modality Alignment. To facilitate effective multi-sensor fusion, we adopt an online contrastive learning strategy to align non-optical features (e.g., SAR) with the optical (RGB) feature space, leveraging the pretrained vision-language projector for optical inputs. Specifically, given a training batch of size $B$, we define the cross-modal contrastive loss as:
$\small\mathcal{L}_{CL}=-\frac{1}{B}\sum_{i=1}^{B}\left(\log\frac{f(x^{o}_{i},x^{s}_{i})}{f(x^{o}_{i},x^{s}_{i})+\sum\limits_{k\neq i}f(x^{o}_{i},x^{s}_{k})}+\log\frac{f(x^{s}_{i},x^{o}_{i})}{f(x^{s}_{i},x^{o}_{i})+\sum\limits_{k\neq i}f(x^{s}_{i},x^{o}_{k})}\right)$ (3)
where $f(\cdot)$ represents the cross-modal cosine similarity operation between paired tokens.
Modality Mutual Attention. After aligning different modalities into a shared semantic space, we further enable the model to dynamically assess the information content of each modality, thereby emphasizing the most informative representations for downstream LLM reasoning.
Specifically, we first introduce a set of learnable queries $Q\in\mathbb{R}^{k\times D}$ to extract neighborhood-aware features from each modality. For example, the representation of SAR is computed via element-wise multiplication: $\hat{X}^{V}_{s}=Q\odot X^{V}_{s}$, where $\hat{X}^{V}_{s}\in\mathbb{R}^{k\times P\times D}$. RGB features are processed similarly to obtain $\hat{X}^{V}_{o}$. We then compute cross-modal importance weights using a dot product:
$\small W_{s}=\text{Softmax}(X^{V}_{o}(\hat{X}^{V}_{s})^{\top}),\quad W_{o}=\text{Softmax}(X^{V}_{s}(\hat{X}^{V}_{o})^{\top})$ (4)
where the weights $W_{s},W_{o}\in\mathbb{R}^{P}$ reflect the relevance of the SAR and RGB tokens in cross-modal contexts, by which we reweight $X^{V}_{s}$ and $X^{V}_{o}$ accordingly and concatenate them into the LLM for joint understanding.
3.4 Training
EarthMind is trained via instruction tuning across diverse supervision formats, including VQA and segmentation tasks. The model learns to generate the target response conditioned on the image tokens, segmentation tokens, and the task instruction. Formally, the generation probability of the next token is defined as:
$\Pr\left(t_{i+1}\mid\underbrace{\texttt{<img>}_{1},\dots,\texttt{<img>}_{P}}_{\mathclap{\scriptstyle\text{image tokens}}},\ \underbrace{\texttt{<seg>}_{1},\dots,\texttt{<seg>}_{Q}}_{\mathclap{\scriptstyle\text{segmentation tokens}}},\ \underbrace{s_{1},\dots,s_{M}}_{\scriptstyle\text{instruction}},\ \underbrace{t_{1},\dots,t_{i}}_{\scriptstyle\text{generated response}};\ \boldsymbol{\Theta}\right)$ (5)
where $\boldsymbol{\Theta}$ denotes the trainable parameters of the LMM. For VQA tasks, we apply standard autoregressive training, minimizing the token-level cross-entropy loss over the ground-truth responses. For segmentation tasks, we use a combination of pixel-wise cross-entropy Loss and Dice Loss. In addition, the KL loss described in Sec. 3.2 is applied to supervise attention-based spatial priors.
Figure 3: Examples of EarthMind-Bench. There are ten tasks to evaluate the multi-sensor fusion ability of LMMs, covering perception and reasoning levels. The LLMs are asked to solve the problem (with the ground-truth answers marked in blue) based on RGB, SAR or RGB-SAR fusion settings.
4 EarthMind-Bench
We observe that existing EO benchmarks lack support for multi-granular understanding and multi-sensor fusion. To address this limitation, we propose EarthMind-Bench, a new benchmark constructed from high-quality RGB–SAR paired data across various public datasets, including OpenEarthMap-SAR [44], DFC2023 Track2 [45], WHU-OPT-SAR [46], MSAW [47], and MultiResSAR [48]. We curate 2,000 samples from their test sets, and design a suite of 10 tasks spanning perception and reasoning, enabling a comprehensive evaluation of LMMs in EO scenarios, as shown in Fig. 3.
Perception. We include six tasks to evaluate both coarse- and fine-grained visual understanding, such as scene classification, object existence, object counting, and image captioning. In particular, we highlight two challenging tasks: 1) Hallucination detection, which requires models to answer some leading questions, and 2) Referring expression segmentation, where models must produce binary masks conditioned on natural language queries.
Reasoning. We organize reasoning tasks into two categories. First, spatial reasoning tasks include spatial relationship, where models infer the relative positions of given objects, and route planning, which requires generating a feasible path from a specified start point to an end point. Second, causal reasoning tasks consist of disaster forecasting and urban development assessment, where models are expected not only to make predictions or judgments, but also to provide reasoning evidence.
All tasks are formatted as either multiple-choice (MC) or open-ended (OE) questions. For MC tasks, we report average accuracy across tasks as the evaluation metric. For OE tasks, we adopt GPT-based scoring to assess the quality of generated responses. All annotations are produced by human experts and have undergone a rigorous verification process, which is described in detail in Appendix D.
5 Experiments
5.1 Implementation Details
EarthMind builds upon InternVL2 [49] and SAM2 [28], adopting a curriculum learning strategy to fine-tune the model across three progressive stages. First, we enhance the instruction-following capability using 1.7M general image-text data, which covers image-level captioning, VQA, region-level object understanding and text-driven segmentation. Second, we introduce 1M EO-specific multimodal data to adapt EarthMind to the remote sensing domain. Third, we utilize our synthesized multi-sensor conversation corpus, and selectively retain examples from earlier stages to mitigate catastrophic forgetting. We fine-tune EarthMind with a learning rate of 4e-5 and a batch size of 2, training only the vision-language projector, the LLM via LoRA [50], and the mask decoder. All experiments are conducted on 8 NVIDIA A100-80G GPUs. More details about training datasets and details can be seen in Appendix F.
5.2 Benchmarks
In addition to our proposed EarthMind-Bench, we further evaluate the effectiveness of EarthMind on several widely-used EO benchmarks, assessing its capability in both multi-granularity and multi-sensor understanding. First, we conduct classification and vision-language evaluations on AID [51], UCMerced [52], RSVQA-HRBEN [53] and the VQA task of VRSBench [35] to perform image-level evaluation. AID is a large-scale aerial image classification dataset consisting of 10,000 images from 30 scene categories, collected via Google Earth. We follow the evaluation protocol of [14], using 20% of the dataset for testing. UCMerced contains 2,100 aerial images from 21 land use classes, commonly used for remote sensing scene classification. RSVQA-HRBEN consists of 47k question-answer pairs, in which we follow [14] to test on the “presence” and “Comparison” categories. The VQA part of VRSBench contains 123k samples covering diverse EO scenarios. Second, for region-level evaluation, we evaluate region-level understanding across region-level captioning and visual grounding task, where DIOR-RSVG [54] and visual grounding task of VRSBench are employed. Third, we conduct referring expression segmentation on the test sets of RefSegRS [55] and RRSIS-D [56] for pixel-level grounding, which include 1,817 and 3,481 test instances, respectively. Last, we involve the test set of BigEarthNet [57], SoSAT-LCZ42 [58], which consists of multispectral data and SAR ship detection dataset [59] to prove our multi-sensor understanding ability.
[…]
Model Size M-Avg O-Avg Scene Class. Object Exist. Halluci. Detect. Object Count. Spatial Relation. Referring Segmen. Image Caption Disaster Forecast. Route Plann. Urban Assess.
Full mark 100 5 100 100 100 100 100 100 5 5 5 5
Evaluation on RGB modality
[…]
GeoChat [14] 7B 41.6 1.78 71.3 51.8 46.8 18.9 19.0 - 1.92 1.73 1.33 2.14
LHRS-bot [15] 7B 47.4 1.84 76.0 58.3 58.3 25.2 19.4 - 2.56 1.75 1.55 1.50
Skysensegpt [37] 7B 45.7 1.45 77.1 56.8 56.7 26.8 11.1 - 1.68 1.40 1.18 1.55
GeoPixel [32] 7B 59.4 2.06 87.2 67.8 73.6 33.5 34.7 46.8 2.80 1.80 1.68 1.95
EarthMind 4B 69.0 2.82 96.5 81.2 81.6 47.3 38.4 54.0 3.35 3.37 2.01 2.55
Evaluation on SAR modality
[…]
GeoChat [14] 7B 40.2 1.45 58.1 49.8 46.8 27.9 18.5 - 1.78 1.65 1.25 1.56
LHRS-bot [15] 7B 41.0 1.71 58.5 56.3 48.5 23.5 18.1 - 1.86 1.70 1.45 1.83
Skysensegpt [37] 7B 39.9 1.55 51.9 53.8 49.2 33.8 10.7 - 1.76 1.70 1.35 1.38
GeoPixel [32] 7B 53.0 1.80 76.8 59.0 65.7 30.5 32.8 35.9 2.08 1.97 1.45 1.68
EarthMind 4B 67.5 2.64 95.4 77.4 74.6 46.8 43.1 53.0 3.10 3.25 1.89 2.30
Evaluation on RGB-SAR Fused modality
[…]
EarthMind 4B 70.6 3.02 97.7 82.4 85.4 47.3 40.3 54.5 3.80 3.37 2.21 2.70
5.3 Main Results
Results on EarthMind-Bench. EarthMind-Bench supports evaluation under three settings: RGB only, SAR only, and RGB–SAR fusion. We compare EarthMind with state-of-the-art EO-specific LMMs and proprietary models such as GPT-4V [1] and GPT-4o [60]. Tab. 2 summarizes the results, from which we highlight three key findings: 1) Fine-grained and open-ended tasks remain challenging for existing LMMs. While coarse-grained tasks like scene classification are largely solved, tasks such as object counting and spatial relationship understanding still exhibit significant performance gaps. In particular, referring segmentation proves difficult for most models due to the lack of pixel-level reasoning ability. 2) Most models generalize poorly to SAR inputs. Performance under SAR-only settings lags behind RGB settings for nearly all models, likely due to training data limitations. In contrast, EarthMind demonstrates strong generalization to SAR inputs, benefiting from multi-sensor training. 3) Effective fusion requires learning cross-modal complementarity, not just stacking modalities. As the only open-source model with explicit fusion capability, EarthMind is compared against GPT-4 variants by feeding both RGB and SAR as pseudo-RGB inputs using the official multi-image interface. Results show that GPT-4 models experience performance degradation compared to RGB-only inputs, particularly on fine-grained tasks like Route Planning, Object Counting, and Spatial Relationship Understanding, where SAR data provides robust structural cues under adverse conditions. In contrast, EarthMind effectively captures cross-modal complementarity, yielding consistent improvements over single-modality inputs.
Table 3: Quantitative performance of EarthMind on public benchmarks. For image-level evaluation, we report accuracy on AID, UC-Merced, RSVQA-HRBEN, and the VQA task of VRSBench. For region-level evaluation, we report CIDEr scores on DIOR-RSVG and Acc@0.5 on the Visual Grounding task of VRSBench. For pixel-level tasks, mean Intersection over Union (mIoU) is reported.
Method Image-level Region-level AID UC RSVQA VRS-VQA RSVG VRS-VG GPT-4o [60] 74.7 88.8 - - - - LLaVA-1.5 [6] 72.0 84.4 63.1 76.4 - 5.1 GeoChat [14] 72.0 84.4 72.3 76.0 30.9 49.8 EarthGPT [17] - - 72.0 - 232.8 - EarthMarker [30] 78.0 86.5 - - 379.3 - EarthDial [31] 88.8 92.4 72.5 - - - EarthMind 97.2 95.0 74.0 78.9 428.2 55.6
Method Pixel-level RRSIS-D RefSegRS LAVT [61] 56.8 47.4 RIS-DMMI [62] 60.3 52.2 Caris [63] 62.1 42.7 RM-SIN [56] 64.2 42.6 CroBIM [64] 64.5 59.8 GeoPixel [32] 67.3 - EarthMind 82.2 62.6
[…]
Method Multispectral BigEarthNet SoSAT-LCZ42 GPT-4o [60] 49.0 15.5 Qwen2.5-VL†[65] 36.2 18.9 EarthDial [31] 69.9 60.7 EarthMind 70.4 58.3
Method SAR Small Medium Large Single Multiple GPT-4o [60] 0.70 0.90 3.20 1.20 0 GeoChat†[14] 2.61 4.92 6.95 2.58 1.40 EarthDial [31] 12.14 26.02 35.56 26.03 6.06 EarthMind 13.58 28.55 36.78 27.45 6.99
Results on Public Benchmarks. We evaluate EarthMind on a series of mainstream EO benchmarks (see Tab. 3). EarthMind consistently delivers strong performance across multiple levels of understanding. On image-level tasks such as scene classification (AID, UC-Merced) and EO visual question answering (RSVQA-HRBEN and VRSBench-VQA), it significantly outperforms previous models, including GPT-4o, despite using only 4B parameters. Moving to region-level tasks, EarthMind achieves a CIDEr score of 428.2 on DIOR-RSVG and 55.6% accuracy on VRSBench visual grounding, outperforming visual-prompt-based methods and highlighting the strength of our region encoder. When it comes to pixel-level benchmarks, EarthMind achieves top results on both RRSIS-D and RefSegRS, surpassing even specialized segmentation models and EO-focused LMMs. Beyond the RGB domain, EarthMind also demonstrates competitive results on multi-sensor understanding, as shown in Tab. 4, demonstrating the strong generalization towards complex EO scenarios.
Figure 4: (Left) Visualization of attention maps between segmentation tokens and image tokens within LLM. The top row is the results of “w/o SAP” and the bottom row is the results of “w/ SAP”, where the attention is reallocated to the region specified by the query. (Right) Comparison of our modality-mutual attention with other baselines under the “Token drop” study.
Table 5: Ablation on Spatial Attention Promoting (Left) and Cross-Modal Fusion (Right)
. Settings RefSegRS RRSIS-D mIoU $\uparrow$ oIoU $\uparrow$ mIoU $\uparrow$ oIoU $\uparrow$ w/o SAP 46.7 86.0 67.5 94.8 w/ SAP (first layer) 46.4 86.0 67.7 95.0 w/ SAP (middle layer) 48.0 86.9 71.8 95.5 w/ SAP (last layer) 47.2 86.5 69.2 95.0 w/ SAP (all layers) 48.4 87.4 72.0 96.0 Settings M-Avg $\uparrow$ mIoU $\uparrow$ w/o Modality Alignment 69.0 53.6 Naive Concatenate 68.4 52.7 Naive Attention 68.6 52.5 Ours 70.6 54.5
5.4 Ablations
EarthMind significantly enhances multi-granularity and multi-sensor EO understanding through the proposed (i) Spatial Attention Prompting (SAP) and (ii) Cross-modal Fusion. We delve into each of these components in the following sections. More discussions can be seen in Appendix G.
Effectiveness of SAP. First, we evaluate the effectiveness of our proposed SAP method. To ensure a fair comparison, we train models from scratch on the training splits of RefSegRS and RRSIS-D, and report results on their corresponding test sets. As shown in Tab. 5, incorporating SAP significantly improves segmentation performance. Second, we visualize the attention maps between the <SEG> token and image tokens across transformer layers within the LLM. As illustrated in Fig. 4, SAP effectively suppresses attention drift and helps reallocate attention toward the anchor regions corresponding to the query. Lastly, we investigate how the performance of SAP varies with the supervision layer. Results show that applying supervision at middle layers yields the strongest performance boost, as these layers balance low-level spatial detail and high-level semantics. Importantly, SAP introduces no additional parameters, making it a lightweight yet effective enhancement for pixel-level grounding within LLMs.
Effectiveness of Cross-modal Fusion. To evaluate the effectiveness of our cross-modal fusion design, we conduct ablations on the multi-sensor setting with three configurations: (1) w/o Modality Alignment; (2) Naive Concatenate: visual tokens from different modalities are directly concatenated and passed to the LLM; (3) Naive Attention: token importance is computed using cosine similarity between paired SAR and optical features. We report both multiple-choice accuracy and referring expression segmentation results on EarthMind-Bench. As shown in Tab. 5, our full model significantly outperforms all baselines, demonstrating the importance of both alignment and adaptive weighting for robust fusion. Furthermore, to assess the informativeness of modality-selected tokens, we compare different token retention strategies. For baselines, we randomly discard a portion of the input tokens before feeding them into the LLM. In contrast, our method selectively drops tokens with lower learned importance weights. As shown in Fig. 4, even under the same retention ratio, EarthMind retains more informative content, leading to better performance. This demonstrates that our method reserves the most complementary information across them.
6 Limitations and Conclusion
Limitations of EarthMind. Training EarthMind demands considerable computational resources due to its use of multiple visual encoders for multi-level understanding. A promising direction is to optimize the architecture via Mixture-of-Experts or knowledge distillation to reduce redundancy. Additionally, a modality-aligned encoder that jointly embeds heterogeneous sensor inputs into a shared semantic space could further improve efficiency.
In this work, we propose EarthMind, a unified framework for multi-granular and multi-sensor understanding in Earth Observation (EO). To enhance pixel-level performance, we introduce Spatial Attention Prompting (SAP), which guides attention reallocation within the LLM. We also design a Cross-modal Fusion mechanism that aligns heterogeneous modalities into a shared space and adaptively selects informative features. To support evaluation, we curate EarthMind-Bench, the first benchmark tailored for multi-sensor fusion in LMMs. Extensive experiments validate the effectiveness and strong generalization of EarthMind across diverse EO tasks.
Appendix A Overview of Appendix
  • B: Discussions and Borader Impact.
  • C: Details of EarthMind.
  • D: Details of EarthMind-Bench.
  • E: Experimental Settings.
  • F: Training Dataset.
  • G: More Ablation Studies.
  • H: More Visualization Results.
Appendix B Discussions and Broader Impact
We discuss the limitation of EarthMind in Sec. 6, here we further discuss the limitation of EarthMind-Bench. Currently, EarthMind-Bench includes only paired optical (RGB) and SAR modalities. To enable a more comprehensive evaluation of LMMs’ multi-sensor fusion capabilities, future extensions should incorporate additional sensing modalities such as multispectral, hyperspectral, and infrared imagery. Moreover, benchmarking diverse fusion scenarios involving more than two modalities would further reflect EO challenges.
Broader Impact. EarthMind offers a broader impact to the large multimodal model (LMM) community in both methodology and application. On the one hand, it demonstrates how vision-language models can be extended to handle multi-granular tasks via the proposed Spatial Attention Prompting mechanism, which reallocates LLM attention to task-relevant regions. This architectural design is not limited to Earth Observation (EO), but can be generalized to other domains requiring fine-grained spatial reasoning, such as medical imaging and autonomous driving.
On the other hand, for the EO community, EarthMind contributes both algorithmically and empirically: it introduces an effective multi-sensor fusion framework and proposes a scalable training pipeline, accompanied by a new benchmark and instruction-tuning dataset. These efforts can benefit various downstream applications in remote sensing, including disaster forecasting, infrastructure monitoring, and route navigation.
Appendix C Details of EarthMind
EarthMind is built upon the InternVL-2 framework [49]. In InternVL-2, each image is divided into multiple patches at pre-defined scales. Each patch is processed by the visual encoder and encoded into 256 tokens. For instance, an image with 4 patches (plus a global image token) yields $(4+1)\times 256$ visual tokens in total. For region-level perception, we adopt the region encoder from GPT4RoI [66], where each detected region is encoded into 256 tokens as well. To support multi-sensor input, non-optical imagery (e.g., SAR or multispectral data) is transformed into a synthetic video-like sequence by stacking pre-processed frames. This sequence is then concatenated before the input query, following the protocol of multi-frame vision-language models.
In addition, we extend the tokenizer by introducing a special “[SEG]” token for pixel-level grounding. The hidden state of the final LLM layer corresponding to the “[SEG]” token serves as the semantic prompt for mask generation. During inference, if the “[SEG]” token is not generated, we interpret it as an indication that the queried object is not present in the image.
Figure 5: Detailed data statistics of the EarthMind-Bench.
Appendix D Details of EarthMind-Bench
Overview. We present a more detailed analysis of EarthMind-Bench in Fig. 5. Subfigures (a–c) illustrate the distribution of task categories, while (d) and (e) report the number of samples per task. Notably, our benchmark includes 438 referring expression segmentation samples paired with corresponding binary masks, highlighting its support for fine-grained pixel-level grounding. We further visualize the word clouds of questions (f) and answers (g), showing that EarthMind-Bench covers a wide variety of object types and semantics, enabling comprehensive evaluation across perception and reasoning tasks.
Data Annotations. Among the 10 subtasks in EarthMind-Bench, annotations for Scene Classification and Referring Expression Segmentation are directly inherited from their original datasets. For the remaining 8 tasks, we rely on human annotations. We recruit 8 domain experts in geoscience, each assigned to annotate a specific task. After the initial annotation phase, all experts cross-validate each other’s work and score the quality of samples. Only high-quality samples with consensus are retained in the final benchmark. To reduce the annotation burden, we first leverage GPT-4o to generate detailed image descriptions, which annotators can reference when constructing question-answer pairs. To ensure consistency and annotation fidelity, we also provide strict task-specific annotation guidelines for all subtasks, which are shown as follows:
  • 1. Object Existence Determine whether a specific object or class exists in the image (e.g., Is there a bridge in the image?”). Guideline: The object should be clearly visible and occupy a non-trivial area. If heavily occluded or ambiguous, mark as Not present.”
  • 2. Hallucination Detection Detect whether the model falsely recognizes objects that do not exist. Guideline: Ask for fine-grained distinctions between semantically similar objects (e.g., “Is there a train or a bus in the image?”). Confirm that the mistaken object does not exist even under partial occlusion.
  • 3. Object Counting Count the number of objects from a given category (e.g., “How many buildings are visible in the image?”). Guideline: Only count objects that are visually distinguishable and not clustered into ambiguous shapes. Accept a small margin of error (±1) in complex scenes.
  • 4. Spatial Relationship Identify relative positions between objects (e.g., What is next to the waterbody?”). Guideline: Use cardinal or contextual spatial relations (e.g., to the left of,” adjacent to,” surrounded by”). Annotators should verify object co-existence and relative proximity.
  • 5. Route Planning Generate or select feasible navigation routes from a start to a target location. Guideline: Ensure the proposed path avoids obstacles (e.g., rivers, buildings), follows terrain constraints (e.g., avoid steep hills), and adheres to logical movement (e.g., roads preferred over fields).
  • 6. Image Captioning Generate descriptive sentences summarizing the content and layout of the image. Guideline: Captions should mention key objects, land types, spatial layout, and human-made structures if visible. Avoid hallucination and be concise yet informative.
  • 7. Disaster Forecasting Assess the likelihood of disaster based on visual evidence (e.g., “Is this area prone to flooding?”). Guideline: Use cues such as terrain (lowlands), proximity to water, lack of infrastructure, or signs of previous disaster impact. Avoid speculative answers.
  • 8. Urban Development Assessment Evaluate whether an area is suitable for urban development. Guideline: Consider factors such as flat terrain, absence of natural barriers, existing road access, and land cover type. Annotators should justify suitability with visual cues.
Evaluation Metrics. For multiple-choice tasks, we evaluate model performance using standard accuracy, measuring the percentage of predictions that exactly match the ground-truth option. For open-ended tasks, inspired by [67], we adopt a GPT-4-based evaluation protocol that assesses the alignment between the model-generated answers and human annotations. As illustrated in Fig. 6, we prompt GPT-4 to score the responses based on semantic similarity and correctness, following a structured rubric to ensure consistent evaluation across tasks.
Figure 6: The illustration of open-ended task evaluation prompt.
Appendix E Details of Experimental Settings
We elaborate on the training and inference details of EarthMind. Specifically, we report the hyperparameters in the fine-tuning stage, as shown in Tab. 6.
Hyperparameter Value
Overall batch size 64
Learning rate 4e-5
LR Scheduler Cosine decay
DeepSpeed ZeRO Stage ZeRO-2
Optimizer Adam
Warmup ratio 0.3
Epoch 1
Weight decay 0
Precision bf16
Table 6: Hyperparameters of EarthMind.
Appendix F Training Dataset
EarthMind is trained on large-scale natural image datasets, including LLaVA-665K [6], 56K referring expression data [68, 69], and 214K grounding conversation generation samples [9]. These datasets cover image-level captioning, VQA, and text-driven segmentation. We also incorporate 724K region-level descriptions from the Osprey dataset [70] to improve region-level understanding capacity. Second, we introduce EO-specific multimodal data to adapt EarthMind to the remote sensing domain. This includes 1M VQA data from EarthGPT [17], 142K EO conversations from VRSBench [35], 31K region-level captions from DIOR-RSVG [54], and 21K referring segmentation samples from RRSIS-D [56] and RefSegRS [55]. Moreover, we involve 500k multi-spectral data, including BigEarthNet [57] and SoSAT-LCZ42 [57]. Third, we utilize our synthesized multi-sensor conversation corpus (20K RGB-SAR paired dialogues) and selectively retain examples from earlier stages to mitigate catastrophic forgetting.
Paired Multi-Sensor Training Data Curation. We construct a high-quality paired RGB–SAR training corpus from five publicly available datasets: OpenEarthMap-SAR [44], DFC2023 Track2 [45], WHU-OPT-SAR [46], MSAW [47], and MultiResSAR [48]. All RGB–SAR pairs are sampled from their training splits to avoid overlap with EarthMind-Bench. To scale data generation, we design an automatic synthetic pipeline based on GPT-4o. Since GPT-4o performs better on RGB imagery than SAR, we use RGB images and their classification labels as input prompts. The prompt is tailored for large vision-language models (VLMs), comprising the image and a list of detected object categories (e.g., buildings, roads, water bodies). The model is first asked to generate a comprehensive caption summarizing the scene and spatial relationships among objects. Then, it is instructed to synthesize 3–10 diverse question–answer pairs. These cover various reasoning types, such as: (1) Object Existence (e.g., “Is there a river in the image?”), (2) Counting (e.g., “How many buildings are visible?”), (3) Spatial Relation (e.g., “What is next to the forest?”), (4) Object Localization (e.g., “Where is the road located?”), and (5) Scene-Level Understanding (e.g., “Is this area suitable for agriculture?”). All outputs are returned in a structured JSON format for seamless integration into our training pipeline. In total, we curate 20k synthetic multi-sensor samples, which serve as paired RGB–SAR instruction data for adaptation of multi-sensor LMMs.
Table 7: Ablations on the non-rgb data processing.
Method Multispectral BigEarthNet SoSAT-LCZ42 Three-band grouping 70.4 58.3 Single-band grouping 71.2 59.2
Method SAR Small Medium Large Single Multiple Zero-padding 12.14 26.02 35.56 26.03 6.06 Channel replication 11.08 25.99 34.38 25.99 4.32
Table 8: Ablation study on Multi-granular joint training.
Method Image-level (Accuracy) Region-level Pixel-level (mIoU) AID UC RSVQA VRS-VQA RSVG VRS-VG RRSIS-D RefSegRS Only trained on Image data 97.0 95.1 73.8 77.8 - - - - Only trained on Region data - - - - 379.6 49.8 - - Only trained on segmentation data - - - - - - 77.6 59.3 EarthMind 97.2 95.0 74.0 78.9 428.2 55.6 82.2 62.6
Table 9: Ablation study on joint multi-sensor training.
Method EarthMind-Bench RGB SAR RGB+SAR Only trained on RGB 68.4 30.1 28.4 Only trained on SAR 45.6 59.8 22.3 trained on paired RGB-SAR 69.0 67.5 70.6
Appendix G More Ablation Studies
Ablation on Multi-Sensor Data Processing. One of the key strengths of EarthMind lies in its ability to handle multi-sensor data beyond standard RGB imagery. To better understand the impact of different preprocessing strategies, we conduct a series of experiments focused on the handling of SAR and multispectral (MS) data. For SAR inputs with fewer than three channels, we compare two strategies: (1) Zero padding, where the missing channels are filled with zeros; (2) Channel replication, where existing channels are duplicated to reach three channels. For multispectral data with more than three bands, we evaluate: (1) Three-band grouping, where every three consecutive bands are grouped to form one RGB-like frame; (2) Single-band grouping, where each band is treated as an individual frame, forming a multi-frame sequence. We evaluate each method under the same training settings and report classification accuracy on the test sets of BigEarthNet (for MS) and SAR Ship Detection (for SAR), as summarized in Tab. 7. Our results show that zero padding outperforms channel replication for SAR data, likely because copying channels introduces redundancy and potential noise. For MS data, single-band grouping slightly improves performance due to better spectral resolution, but incurs substantial computational overhead due to the increased number of tokens. Considering both effectiveness and efficiency, we adopt zero padding for SAR and three-band grouping for MS as our default configuration. In the future, we will consider involve token reduction techniques [71] to reduce overhead cost.
Ablation on Joint Multi-Granular and Multi-Sensor Training. EarthMind is designed to be jointly trained on both multi-granular and multi-sensor data. To validate the effectiveness of this unified training paradigm, we conduct ablation studies along two axes: (1) Multi-Granular Training. We compare two settings: (i) Joint training with image-level, region-level, and pixel-level data. (ii) Independent training for each granularity using the same total amount of data. (2) Multi-Sensor Training. We also compare: (i) Joint training with all available sensor modalities (e.g., RGB, SAR, MS). (ii) Independent training with each modality separately. As shown in Tab. 8 and Tab. 9, the multi-granular co-training strategy consistently outperforms independently trained counterparts, especially on pixel-level tasks. This suggests that high-level semantic supervision (e.g., image-level QA) can improve fine-grained understanding through shared representation learning. Similarly, multi-sensor co-training improves generalization across modalities. Notably, the performance under SAR-only evaluation is significantly enhanced by leveraging complementary information learned from RGB and MS data during joint training. This highlights EarthMind’s ability to exploit cross-modal synergy in both perception and reasoning tasks.
Ablation on Curriculum Training Strategy. EarthMind adopts a curriculum learning strategy that gradually adapts the model from general vision-language data to remote sensing tasks. Specifically, the training proceeds in three stages: (1) pretraining on large-scale natural image VQA and captioning datasets, (2) domain adaptation using EO-specific RGB data, and (3) fine-tuning on multi-sensor data (e.g., SAR). To evaluate the effectiveness of this curriculum, we conduct ablation studies by removing or reordering training stages. As shown in Tab. 10 and Tab. 11, pretraining on general image-language data provides strong foundational capabilities, leading to significant performance improvements on EO tasks, particularly in pixel-level segmentation. Furthermore, RGB-domain training not only enhances performance on RGB inputs but also boosts multi-sensor fusion results, demonstrating its role as an effective bridge between general vision and SAR-specific domains.
Table 10: Ablation study on the first stage of curriculum training.
Method Image-level (Accuracy) Region-level Pixel-level (mIoU) AID UC RSVQA VRS-VQA RSVG VRS-VG RRSIS-D RefSegRS w/o pretraining 96.5 94.8 73.2 77.6 406.7 49.8 75.4 55.3 with pretraining 97.2 95.0 74.0 78.9 428.2 55.6 82.2 62.6
Table 11: Ablation study on the second stage of curriculum learning.
Method EarthMind-Bench RGB SAR RGB+SAR w/o pretraining on RGB EO data 67.5 64.3 68.9 with pretraining on RGB EO data 69.0 67.5 70.6
Appendix H More Visualization Results
More visualization results of EarthMind can be seen in Fig. 7.
Figure 7: More visualization of EarthMind.